### v1

# Leveraging Machine Learning for High-Dimensional Option Pricing within the Uncertain Volatility Model

Abstract
This paper explores the application of Machine Learning techniques for pricing high-dimensional options within the framework of the Uncertain Volatility Model (UVM). The UVM is a robust framework that accounts for the inherent unpredictability of market volatility by setting upper and lower bounds on volatility and the correlation among underlying assets. By leveraging historical data and extreme values of estimated volatilities and correlations, the model establishes a confidence interval for future volatility and correlations, thus providing a more realistic approach to option pricing. By integrating advanced Machine Learning algorithms, we aim to enhance the accuracy and efficiency of option pricing under the UVM, especially when the option price depends on a large number of variables, such as in basket or path-dependent options. Our approach evolves backward in time, dynamically selecting at each time step the most expensive volatility and correlation for each market state. Specifically, it identifies the particular values of volatility and correlation that maximize the expected option value at the next time step. This is achieved through the use of Gaussian Process regression, the computation of expectations via a single step of a multidimensional tree and the Sequential Quadratic Programming optimization algorithm. The numerical results demonstrate that the proposed approach can significantly improve the precision of option pricing and risk management strategies compared with methods already in the literature, particularly in high-dimensional contexts.
Keywords: UVM, option pricing, GPR, Tree method, SQP.
1 Introduction
The Uncertain Volatility Model represents a significant advancement in financial modeling, specifically for the pricing and hedging of derivative securities in markets characterized by uncertain volatility. Developed initially by Avellaneda et al. [1], the UVM addresses the limitations of traditional models like the Black-Scholes, which assume constant volatility over the life of an option. Instead, the UVM assumes that volatility fluctuates within a known range, providing a more realistic framework for financial markets where volatility is inherently unpredictable.
The core idea of the UVM is to set upper and lower bounds on volatility, denoted as $\sigma^{\min}$ and $\sigma^{\max}$. These bounds can be inferred from historical data or extreme values of implied volatilities from liquid options. The model uses these bounds to establish a confidence interval within which the future volatility is expected to lie. This approach acknowledges the inherent uncertainty in predicting future volatility and provides a robust framework for pricing options under this uncertainty [1, 8].
One of the seminal contributions of Avellaneda et al. [1] was the derivation of the Black-Scholes-Barenblatt (BSB) equation, a non-linear partial differential equation (PDE) that extends the classical Black-Scholes equation by incorporating the variable volatility bounds. This PDE dynamically selects the "pricing" volatility from the two extreme values based on the convexity of the value function. For instance, if the second derivative (Gamma) of the option price is positive, the model uses $\sigma^{\max}$, otherwise, it uses $\sigma^{\min}$.
The UVM framework is particularly effective for superhedging strategies, where traders aim to hedge against the worst-case scenarios of volatility. By ensuring that the option pricing incorporates the highest potential volatility, traders can guarantee non-negative profit and loss outcomes even when actual realized volatility deviates significantly from the expected values. This is especially useful in markets with high volatility or during periods of market stress, as pointed out by Martini and Jacquier [8].
In addition to its theoretical foundations, the UVM has practical implications for risk management. Pooley et al. [9] provided a detailed numerical analysis of the UVM, highlighting its advantages in constructing efficient hedges and managing derivatives positions in environments with uncertain volatility. Their work emphasized the importance of diversification in managing volatility risk and provided algorithms for implementing the UVM in real-world trading systems.
Several extensions and practical applications of the UVM have been explored in subsequent research. For instance, Windcliff et al. [13] use methods based on partial differential equations (PDEs) to evaluate cliquet options in the context of the uncertain volatility model. The paper explores how to solve the PDEs associated with these financial instruments by comparing different numerical and grid construction techniques. It is shown that the PDE-based approach is effective in capturing the complexities of the UVM model, providing a more accurate evaluation of cliquet options than other traditional methods. Martini and Jacquier [8] elaborated on the superhedging strategies within the UVM framework, discussing the implications for derivative desks where risk aversion is a critical factor. They demonstrated that the UVM could be applied to a range of option types, including those with non-convex payoffs, by employing more sophisticated numerical methods such as trinomial trees and finite-difference schemes.
More recently, Guyon and Labordère [6] introduce two innovative Monte Carlo methods to price financial derivatives under the UVM. These methods address the challenges posed by high-dimensionality, where traditional finite-difference schemes fall short. Moreover, the paper considers a path-dependent option, which is particularly useful in the context of the UVM. This is beneficial because path-dependent options can capture the effects of the asset’s historical prices, providing a more accurate valuation under uncertainty, and aligning well with the UVM’s focus on modeling worst-case scenarios. Despite one may appreciate the innovative approach presented in the paper, it is worth noting that the numerical results are occasionally suboptimal, and the authors do not extend their numerical analysis to derivatives with more than two underlyings.
In this paper, we propose an innovative approach to the problem of derivative evaluation within the UVM. Specifically, the proposed method is based on the GPR-Tree pricing algorithm, which we introduced in a previous paper [5] for the valuation of options dependent on multiple underlyings. Consequently, we refer to this method as GTU, which stands for GPR-Tree method for UVM. Our algorithm involves calculating the price in the UVM by working backward in time and determining, for selected market states, the maximum price by leveraging an enhanced tree method and a suitable optimization algorithm. This calculation accounts for varying volatilities and correlations between the underlyings. The obtained prices for specific typical market states are then generalized to all possible market conditions by using Gaussian Process Regression (GPR), a powerful machine learning technique for multidimensional regressions from scattered data.
The proposed method not only serves as a viable alternative to existing methods for handling problems in lower dimensions but also demonstrates exceptional capability in valuing derivatives in high-dimensional settings, including scenarios with dozens of underlyings. To the best of our knowledge, this is the first instance where such a high-dimensional problem has been addressed effectively. Moreover, the proposed method is particularly effective for problems where the high dimension arises not from the number of underlyings, but from the impact of the entire process history on the derivative’s final value. This is especially relevant for path-dependent options, which our algorithm can evaluate proficiently.
The paper includes the results of several numerical experiments, which show that the GTU algorithm is reliable and fast. In particular, we observe that the results returned by GTU are very stable and very close to the benchmark prices when these are available.
The remainder of the paper is organized as follows. Section 2 introduces the UVM and the pricing principles relevant to this context. Section 3 reviews the fundamentals of the GPR-Tree algorithm. Section 4 demonstrates the evolution of the GPR-Tree algorithm into the proposed GTU method. Section 5 provides detailed numerical results from key experiments. Finally, Section 6 concludes the paper.
2 The Multidimensional Uncertain Volatility Model
Let $\mathbf{S}=(\mathbf{S}_{t})_{t\in[0,T]}$ denote the $d$-dimensional underlying process, which evolves randomly according to the multidimensional Black-Scholes model. Under the risk-neutral probability measure $\mathbb{Q}$, the dynamics of this model are given by the stochastic differential equation:
$dS_{t}^{i}=\left(r-\eta_{i}\right)S_{t}^{i}\,dt+\sigma_{i}S_{t}^{i}\,dW_{t}^{i},\quad i=1,\ldots,d,$ (2.1)
where $\mathbf{S}_{0}=(s_{0}^{1},\dots,s_{0}^{d})\in\mathbb{R}_{+}^{d}$ is the initial spot price vector, $r$ is the constant risk-free interest rate, $\mathbf{\eta}=(\eta_{1},\dots,\eta_{d})^{\top}$ is the vector of dividend yields, $\mathbf{\sigma}=(\sigma_{1},\dots,\sigma_{d})^{\top}$ is the vector of volatilities, $\mathbf{W}$ is a $d$-dimensional correlated Wiener processes, and $\rho_{i,j}$ is the instantaneous correlation coefficient between $W^{i}$ and $W^{j}$. We also term $\left\{\mathcal{F}_{t}\right\}_{t\geq 0}$ the natural filtration generated by the $\mathbf{W}$.
Equation (2.1) can be rewritten as
$dS_{t}^{i}=S_{t}^{i}(\left(r-\eta_{i}\right)dt+\sigma_{i}\Sigma_{i}d\mathbf{B}_{t}),$ (2.2)
where $\mathbf{B}$ is a $d$-dimensional uncorrelated Wiener processes and $\Sigma_{i}$ is the $i$-th row of the matrix
$\Sigma=\Sigma\left(\rho_{1,2},\dots,\rho_{d-1,d}\right)$ (2.3)
defined as a square root of the correlation matrix $\Gamma$, given by
$\Gamma=\Gamma\left(\rho_{1,2},\dots,\rho_{d-1,d}\right)=\begin{pmatrix}1&\rho_{1,2}&\ldots&\rho_{1,d}\\ \rho_{1,2}&1&\ddots&\vdots\\ \vdots&\ddots&\ddots&\rho_{d-1,d}\\ \rho_{1,d}&\ldots&\rho_{d-1,d}&1\end{pmatrix}.$ (2.4)
Moreover, one can write
$\mathbf{S}_{t}=\mathbf{S}_{0}\circ\exp\left(\left(r-\mathbf{\eta}-\frac{\sigma^{2}}{2}\right)\Delta t+\sigma\circ\left(\Sigma\mathbf{B}_{t}\right)\right)$ (2.5)
with $\mathbf{v}\circ\mathbf{w}$ the Hadamard product between vectors $\mathbf{v}$ and $\mathbf{w}$, that is the element-wise product, and the exponential function in (2.5) is applyed to each component of the vector $\left(r-\frac{\sigma^{2}}{2}\right)\Delta t+\sigma\circ\left(\Sigma\mathbf{B}_{t}\right)$.
Instead of assuming a single constant volatility, the UVM considers a range of possible volatilities. Specifically, the vector of volatilities $\mathbf{\sigma}$ is not known, but it can change over time, as a function of the time and of the underlyings, so we write $\mathbf{\sigma}=\sigma\left(t,\mathbf{S}_{t}\right)=\left(\sigma_{1}\left(t,\mathbf{S}_{t}\right),\dots,\sigma_{d}\left(t,\mathbf{S}_{t}\right)\right)^{\top}$. Moreover, the explicit form of $\sigma\left(t,\mathbf{S}_{t}\right)$ is not specified in the model, but we only know that, for each $i=1,\dots,d$ there exist two bounds for $\sigma_{i}$, that is
$\sigma_{i}^{\min}\leq\sigma_{i}\left(t,\mathbf{S}_{t}\right)\leq\sigma_{i}^{\max}.$ (2.6)
Such a degree of uncertainty can be extended to correlations between Wiener processes. Specifically $\rho_{i,j}=\rho_{i,j}\left(t,\mathbf{S}_{t}\right)$ may change over time based on the underlying values, but the particular law is not a model datum. Instead, we only know that, for each couple of values $i=1,\dots,d-1$ and $j=i+1,\dots,d$, there exist two bounds for $\rho_{i,j}$, that is
$\rho_{i,j}^{\min}\leq\rho_{i,j}\left(t,\mathbf{S}_{t}\right)\leq\rho_{i,j}^{\max}.$ (2.7)
To summarise, in the UVM, the dynamics of each underlying $S^{i}$ can be written down as
$dS_{t}^{i}=\left(r-\eta_{i}\right)S_{t}^{i}\,dt+\sigma_{i,t}\left(t,\mathbf{S}_{t}\right)S_{t}^{i}\,dW_{t}^{i},\quad i=1,\ldots,d,$ (2.8)
with $\sigma_{i,t}\left(t,\mathbf{S}_{t}\right):\left[0,T\right]\times\mathbb{R_{+}}^{d}\rightarrow\left[\sigma_{i}^{\min},\sigma_{i}^{\max}\right]$ a function. For every couple of indices $i=1,\dots,d-1$ and $j=i+1,\dots,d$ we have
$d\left\langle W_{t}^{i},W_{t}^{j}\right\rangle=\rho_{i,j}\left(t,\mathbf{S}_{t}\right),$
with $\rho_{i,j}\left(t,\mathbf{S}_{t}\right):\left[0,T\right]\times\mathbb{R_{+}}^{d}\rightarrow\left[\rho_{i,j}^{\min},\rho_{i,j}^{\max}\right]$ a function. Finally, in order to preserve the interpretability of the $\rho_{i,j}$ coefficients as correlations, we require the matrix $\Gamma\left(t,\mathbf{S}_{t}\right)=\Gamma\left(\rho_{1,2}\left(t,\mathbf{S}_{t}\right),\dots,\rho_{d-1,d}\left(t,\mathbf{S}_{t}\right)\right)$ to be positive semidefinite, which we denote with
$\Gamma\left(t,\mathbf{S}_{t}\right)\succeq 0.$ (2.9)
The price of a financial derivative in the UVM can be formulated as a stochastic control problem. To this aim, we denote with
$P=P\left(t,\mathbf{S}_{t}\right)=\left(\sigma_{1}\left(t,\mathbf{S}_{t}\right),\dots,\sigma_{d,t}\left(t,\mathbf{S}_{t}\right),\rho_{1,2}\left(t,\mathbf{S}_{t}\right),\dots,\rho_{d-1,d}\left(t,\mathbf{S}_{t}\right)\right)^{\top}$
the vector of all uncertain model parameters as a function of $t$ and $\mathbf{S}_{t}$, and $\mathcal{A}$ the set of all admissible values $P$, that is, which satisfy relations (2.6), (2.7), and (2.9).
In the UVM, the pricing of derivatives is approached as an optimization problem due to the uncertainty in the volatility and correlation of the underlying asset. This model aims to find the worst-case scenario within this range to ensure robust and conservative pricing. Specifically, we seek to maximize the expected payoff of the derivative over all admissible scenarios for volatilities and correlation, that is over all $P$ in $\mathcal{A}$. Let $V\left(t,\mathbf{S}_{t}\right)$ represent the value of the derivative at time $t$ and asset price $\mathbf{S}_{t}$ in the UVM. The pricing problem is then given by:
$V\left(t,\mathbf{S}_{t}\right)=\sup_{P\in\mathcal{A}}\mathbb{E}\left[e^{-r(T-t)}\Psi\left(\mathbf{S}_{T}\right)\mid\mathbf{S}_{t}\right],$ (2.10)
where $\Psi$ is the payoff function at maturity $T$, the dynamics of $\mathbf{S}_{t}$ is given by (2.8) for $P\in\mathcal{A}$, and $\mathbb{E}$ denotes the expectation under the risk-neutral measure $\mathbb{Q}$.
This optimization ensures that the derivative’s price accounts for the most adverse volatility-correlation scenario, providing a conservative estimate that is robust to volatility fluctuations. By solving this maximization problem, we obtain a price that safeguards against the uncertainty in the market, making it a reliable valuation method in uncertain volatility conditions.
3 The GPR-Tree Method
The proposed GTU method is based on the GPR-Tree algorithm. Here we give some details about this latter algorithm and we refer the interested reader for further details to [5].
The GPR-Tree algorithm exploits Gaussian Gaussian Process Regression and a multimensional binomial tree to compute the price of multidimensional American option.
GPR is a non-parametric kernel-based probabilistic model used for regression tasks. It is particularly useful for making predictions about complex datasets in high dimensions. GPR models the observations as samples from a Gaussian process, which provides a probabilistic framework for prediction.
Let $X=\left\{\mathbf{x}_{p},p=1,\dots,P\right\}\subset\mathbb{R}^{d}$ be a set of $d$-dimensional predictors and $Y=\left\{y_{p},p=1,\dots,P\right\}\subset\mathbb{R}$ be the corresponding set of scalar outputs. These observations are modeled as:
$y_{p}=f(\mathbf{x}_{p})+\epsilon_{p},$ (3.1)
where $f$ is a Gaussian process and $\epsilon_{p}$ is Gaussian noise with variance $\sigma^{2}$.
A Gaussian process $f(\mathbf{x})\sim\mathcal{GP}(m(\mathbf{x}),k(\mathbf{x},\mathbf{x}^{\prime}))$ is fully specified by its mean function $m(\mathbf{x})$ and covariance function $k(\mathbf{x},\mathbf{x}^{\prime})$. The mean function is often assumed to be zero, $m(\mathbf{x})=0$, and the covariance function (kernel) defines the relationship between different points in the input space. A common choice for the kernel is the Matern 3/2 (M3/2) kernel, given by:
$k_{\text{M3/2}}(\mathbf{x},\mathbf{x}^{\prime})=\sigma_{f}^{2}\left(1+\frac{\sqrt{3}}{\sigma_{l}}\|\mathbf{x}-\mathbf{x}^{\prime}\|_{2}\right)\exp\left(-\frac{\sqrt{3}}{\sigma_{l}}\|\mathbf{x}-\mathbf{x}^{\prime}\|_{2}\right),$ (3.2)
where $\sigma_{f}^{2}$ is the signal variance and $\sigma_{l}$ is the length scale.
Given a new input $\mathbf{x}^{*}\in\mathbb{R}^{d}$, the predicted mean and variance of the corresponding output are given by:
$\mu_{*}=K(\mathbf{x}^{*},X)[K(X,X)+\sigma_{n}^{2}I_{n}]^{-1}Y,$ (3.3)
$\sigma_{*}^{2}=K(\mathbf{x}^{*},\mathbf{x}^{*})-K(\mathbf{x}^{*},X)[K(X,X)+\sigma_{n}^{2}I_{n}]^{-1}K(X,\mathbf{x}^{*}),$ (3.4)
where $K(X,X)$ with $K\left(X,X\right)_{i,j}=k\left(\mathbf{x}_{i},\mathbf{x}_{j}\right)$ is the covariance matrix computed using the kernel function for all pairs of training inputs, $\sigma_{n}^{2}$ is the noise volatility and $I_{n}$ is the $n\times n$ identity matrix. The parameters $\sigma_{f},\sigma_{l}$ and $\sigma_{n}$ are referred to as hyperparameters and are estimated through log-likelihood maximization. The value $f(\mathbf{x})$ is predicted by $\mu_{*}$, while $\sigma_{*}^{2}$ can be exploited to create confidence intervals for such a prediction. Therefore, GPR is an effective tool for performing multidimensional regressions on sparse sets of input data points.
The GPR-Tree method proceeds by approximating the price of an American option with a Bermudan option on the same basket. At each exercise date $t_{n}$, the option value is computed as the maximum between the exercise value and the continuation value, which is approximated using GPR.
Let us consider a $d$-dimensional Black-Scholes model, describer by equations (2.1). Let $N$ be the number of time steps, $\Delta t=T/N$, and $t_{n}=n\Delta t$. At any exercise date $t_{n}$, the value of the option is:
$V(t_{n},\mathbf{S}_{t_{n}})=\max\left(\Phi(S_{t_{n}}),C(t_{n},\mathbf{S}_{t_{n}})\right),$ (3.5)
where $C(t_{n},S_{t_{n}})$ denotes the continuation value given by:
$C(t_{n},\mathbf{S}_{t_{n}})=\mathbb{E}\left[e^{-r\Delta t}V(t_{n+1},\mathbf{S}_{t_{n+1}})\mid\mathbf{S}_{t_{n}}\right].$ (3.6)
To approximate the continuation value $C(t_{n},S_{t_{n}})$, the GPR-Tree method uses a set $X_{n}$ of predetermined points, whose elements represent certain possible values for the underlyings $\mathbf{S}$:
$X^{n}=\left\{\mathbf{x}^{n,p}=\left(x_{1}^{n,p},\dots,x_{d}^{n,p}\right)^{\top},p=1,\dots,P\right\}\subset\left]0,+\infty\right[^{d}.$ (3.7)
The elements of $X^{n}$ are obtained through a quasi-random simulation of $\mathbf{S}_{T}$ based on the Halton sequence, which allows one to cover the region of interest, avoiding to leave some uncovered areas and to create useless clusters of points (other low discrepancy sequences may be considered, such as Sobol’s one). Specifically,
$\mathbf{x}_{i}^{n,p}=\mathbf{S}_{0}^{i}\exp\left(\left(r-\eta_{i}-\frac{\sigma_{i}^{2}}{2}\right)t_{n}+\sigma_{i}\sqrt{t_{n}}\Sigma_{i}\Phi^{-1}\left(\mathbf{h}^{q}\right)\right),$ (3.8)
with $\mathbf{h}^{q}$ the $q$-th point (column vector) of the Halton sequence in $\left[0,1\right]^{d}$ and $\Phi^{-1}$ stands for the cumulative distribution function of a standarn Gaussian Variable, then applied to each component of $\mathbf{h}^{q}$. The GPR-Tree method assesses the continuation value through one step of the binomial tree proposed by Ekval [4]. For each point $\mathbf{x}^{n,p}$, we consider a set $\tilde{X}^{n,p}$ of $M=2^{d}$ possible values for $\mathbf{S}_{t_{n+1}}\mid\mathbf{S}_{t_{n}}=\mathbf{x}^{n,p}$, which are computed as follows:
$\tilde{\mathbf{x}}_{i}^{n,p,m}=\mathbf{x}_{i}^{n,p}\exp\left(\left(r-\eta_{i}-\frac{\sigma_{i}^{2}}{2}\right)\Delta t+\sigma_{i}\sqrt{\Delta t}\Sigma_{i}\mathbf{G}_{m}\right),$ (3.9)
being $\mathbf{G}_{m}$ the $m$-th point of the space $\left\{-1,+1\right\}^{d}$. It is worth noticing that, as pointed out in [4], the elements of $\tilde{X}^{n,p}$ are equally likely and this simplifies the evaluation of the expected value to the computation of the arithmetic mean of the future values. Generally speaking, given $\tilde{V}_{n+1}\left(\cdot\right)$ a suitable approximation of the option value $V\left(t_{n+1},\cdot\right)$, the price function at time $t_{n}$ for $\mathbf{S}_{t_{n}}=\mathbf{x}^{n,p}$ is approximated by
$V_{n}^{Tree}\left(\mathbf{x}^{n,p}\right)=\max\left(\Psi\left(\mathbf{x}^{n,p}\right),\frac{e^{-r\Delta t}}{2^{d}}\sum_{m=1}^{2^{d}}\tilde{V}_{n+1}\left(\tilde{\mathbf{x}}^{n,p,m}\right)\right).$ (3.10)
In order to reduce the computational complexity of (3.10) when $d$ is high, see Remark 2 in Section 4. Then, GPR is employed to achive an approximation of $V\left(t_{n},\cdot\right)$ form the set $X_{n}$ of the predictors and the set $\left\{V_{n}^{Tree}\left(\mathbf{x}^{n,p}\right),p=1,\dots,P\right\}$ of the corresponding outputs. This procedure is repeated backward in time untile time $t=0$ is reached.
4 The GTU Method
Avellaneda et al. [1] propose solving the dynamic optimization problem (2.10) by dynamic programming, and here we take the same backward approach. First of all, we consider a finite number of time steps $N$, a time interval $\Delta t=T/N$ and a discrete set of times $\left\{t_{n}=n\cdot\Delta t,n=0,\dots,N\right\}$. The option price at contract inception $V\left(0,\mathbf{S}_{0}\right)$ can be computed by backward induction, starting from $t_{N}=T$. In particular, the terminal condition is given by
$V\left(T,\mathbf{S}_{T}\right)=\Psi\left(\mathbf{S}_{T}\right).$
Now, let $\tilde{V}_{n+1}\left(\cdot\right)$ be a suitable approximation of $V\left(t_{n+1},\cdot\right)$ and let us suppose $\mathbf{x}$ to be a fixed point of $\mathbb{R}^{d}.$ Then one can define $\tilde{V}_{n}\left(\cdot\right)$ be a suitable approximation of $V\left(t_{n},\cdot\right)$ at $\mathbf{x}$ as
$\tilde{V}_{n}\left(\mathbf{x}\right)=\sup_{P\in\mathcal{A}}\mathbb{E}\left[e^{-r\Delta t}\tilde{V}_{n+1}\left(\mathbf{S}_{t_{n+1}}\right)\mid\mathbf{S}_{t}=\mathbf{x}\right].$ (4.1)
Formally, the model parameters $P$ can vary between $t_{n}$ and $t_{n+1}$, however, following Avellaneda et al. [1], the optimal solution to the problem (4.1) can be approximated by considering a constant function with respect both time and underlying values. So, one considers the problem
$\tilde{V}_{n}\left(\mathbf{x}\right)=\sup_{P\in\mathcal{B}}\mathbb{E}\left[e^{-r\Delta t}\tilde{V}_{n+1}\left(\mathbf{S}_{t_{n+1}}\right)\mid\mathbf{S}_{t_{n}}=\mathbf{x}\right],$ (4.2)
with $\mathcal{B}$ the set of parameter functions which are constant. This is particularly interesting because, a constant function is identified by the value it takes. For a particular $\mathbf{x}$, we can consider the following constant functions
$\sigma_{1}\left(t,\mathbf{S}_{t}\right)=\hat{\sigma}_{1,\mathbf{x}},\dots,\sigma_{d,t}\left(t,\mathbf{S}_{t}\right)=\hat{\sigma}_{d,\mathbf{x}},\rho_{1,2}\left(t,\mathbf{S}_{t}\right)=\hat{\rho}_{1,2,\mathbf{x}},\dots,\rho_{d-1,d}\left(t,\mathbf{S}_{t}\right)=\hat{\rho}_{d-1,d,\mathbf{x}}$
a total of $d\left(d+1\right)/2$ parameters and we set
$D=\left[\sigma_{1}^{\min},\sigma_{1}^{\max}\right]\times\dots\times\left[\sigma_{d}^{\min},\sigma_{d}^{\max}\right]\times\left[\rho_{1,2}^{\min},\rho_{1,2}^{\max}\right]\times\dots\times\left[\rho_{d-1,d}^{\min},\rho_{d-1,d}^{\max}\right]$
$C=\left\{c_{\mathbf{x}}=\left(\hat{\sigma}_{1,\mathbf{x}},\dots,\hat{\sigma}_{d,\mathbf{x}},\hat{\rho}_{1,2,\mathbf{x}},\dots,\hat{\rho}_{d-1,d,\mathbf{x}}\right)\in D\ s.t.\ \Gamma\left(\hat{\rho}_{1,2,\mathbf{x}},\dots,\hat{\rho}_{d-1,d,\mathbf{x}}\right)\succeq 0\right\}$
$\tilde{V}_{n}\left(\mathbf{x}\right)=\sup_{c_{\mathbf{x}}\in C}\mathbb{E}\left[e^{-r\Delta t}\tilde{V}_{n+1}\left(\mathbf{x}\cdot\exp\left(\left(r-\eta-\frac{\sigma^{2}}{2}\right)\Delta t+\left(\hat{\sigma}_{1,\mathbf{x}},\dots,\hat{\sigma}_{d,\mathbf{x}}\right)^{\top}\circ\left(\Sigma\left(\hat{\rho}_{1,2,\mathbf{x}},\dots,\hat{\rho}_{d-1,d,\mathbf{x}}\right)\left(\mathbf{B}_{t_{n+1}}-\mathbf{B}_{t_{n}}\right)\right)\right)\right)\right].$ (4.3)
To address problem (4.3), we must overcome two key challenges: determining the expected value and maximizing such an expected value.
Our proposed method, the GTU algorithm, involves calculating the expected value through a single tree step, following the same approach as described in the previous Section 3. Note that this choice is particularly effective in that, once a set of parameters $c_{\mathbf{x}}$ is fixed, this calculation can be handeld as for the standard multidimensional Blasck-Scholes model:
$\frac{1}{2^{d}}\sum_{m=1}^{2^{d}}e^{-r\Delta t}\tilde{V}_{n+1}\left(\mathbf{x}\cdot\exp\left(\left(r-\eta-\frac{\sigma^{2}}{2}\right)\Delta t+\left(\hat{\sigma}_{1,\mathbf{x}},\dots,\hat{\sigma}_{d,\mathbf{x}}\right)^{\top}\circ\left(\Sigma\left(\hat{\rho}_{1,2,\mathbf{x}},\dots,\hat{\rho}_{d-1,d,\mathbf{x}}\right)\mathbf{G}_{m}\right)\sqrt{\Delta t}\right)\right)$ (4.4)
with $\mathbf{G}_{m}$ the $m$-th point of the space $\left\{-1,+1\right\}^{d}$
Once we are able to compute the expected value for any $c_{\mathbf{x}}\in C$ fixed, we can tackle the optimization problem: this is a multidimensional optimisation problem with inequality constraints that are both linear (the limits for the parameters) and non-linear (the positivity condition of the $\Gamma$ matrix). This is done by means of the Sequential Quadratic Programming (SQP) algorithm, a powerful and widely used method for solving nonlinear optimization problems, particularly those with constraints. SQP methods solve a sequence of optimization subproblems, each of which approximates the original nonlinear problem by a quadratic programming problem. Drawing on the research by Biggs [2], Han [7], and Powell [10, 11], this approach enables a close simulation of Newton’s method for constrained optimization, similar to its application in unconstrained optimization. Sequential Quadratic Programming methods are considered the pinnacle of nonlinear programming techniques. For instance, Schittkowski [12] developed and evaluated a version that surpasses all other tested methods in efficiency, accuracy, and success rate across a wide array of test problems. For this reason, we have propose to use the SQP algorithm to menage problem (4.3).
The procedure described above allows us to calculate the UVM price for a particular value $\mathbf{x}$ of $\mathbf{S}_{t}$, at time $t_{n}$, knowing (an approximation of) the price function $\tilde{V}_{n+1}$ at time $t_{n+1}$.
In order to actively implement this procedure so as to calculate the price at initial time $t=0$, one problem remains to be solved: obtaining an analytical expression of $\tilde{V}_{n}$ from observations $\tilde{V}_{n}\left(\mathbf{x}\right)$ of the same function at certain points. To do this, we use GPR regression. As described in Section 3, at each time-step $n=1,\dots,N$, we first create a mesh of points $X^{n}$ as in (3.7), but in this case the points $\mathbf{x}^{n,p}$ are define through
$\mathbf{x}_{i}^{n,p}=\mathbf{S}_{0}^{i}\exp\left(\left(r-\frac{\left(\sigma_{i}^{\text{avg}}\right)^{2}}{2}\right)t_{n}+\left(\sigma_{i}^{\text{avg}}\right)^{2}\sqrt{t_{n}}\Sigma_{i}^{\text{avg}}\Phi^{-1}\left(\mathbf{h}^{q}\right)\right),$ (4.5)
where
$\sigma_{i}^{\text{avg}}=\frac{\sigma_{i}^{\text{$\min$}}+\sigma_{i}^{\text{$\max$}}}{2},$ (4.6)
and $\Sigma_{i}^{\text{avg}}$ is the square root of the correlation matrix $\Gamma^{\text{avg}}$, given by
$\Gamma^{\text{avg}}=\Gamma\left(\rho_{1,2}^{\text{avg}},\dots,\rho_{d-1,d}^{\text{avg}}\right),$ (4.7)
with
$\rho_{i,j}^{\text{avg}}=\frac{\rho_{1,2}^{\min}+\rho_{1,2}^{\max}}{2}.$ (4.8)
Then, we estimate the price function by using the approximated solution of (4.3), and then use GPR to estimate the function $\tilde{V}_{n}$ outside this set of points. Moving backwards, it is finally possible to calculate an estimate of the option price at time $t=0$, by considering $X_{0}=\left\{\mathbf{S}_{0}\right\}.$
Remark 1.
Unlike neural networks, GPR requires few observations to produce good estimates of the function to be regressed, typically a few thousand. This feature combines well with the fact that, for each time-step $n$ and for each point in $X_{n}$, an optimization problem must be solved.
Remark 2.
The summation that appears in the formula (4.4) contains a number of addends that grows exponentially with respect to the dimension $d$. For large dimensions, one therefore has to consider the arithmetic mean of a large number of addends. A possible approach to limit this quantity is to consider only a certain number of addends $M\leq 2^{d}$ sampled randomly among the $2^{d}$ addends, and make $M$ tend towards $2^{d}$. Furthermore, to improve the predictive power of this reduced sample, assuming $M$ a even number, we sample the $G_{m}$ values needed to calculate the addends in (4.3) by antithetic variables. This means that if we include in the summation the value that is obtained from a certain $G_{m}$ value, we also include that which is generated from $-G_{m}$. This expedient improves the convergence of the procedure.
As demonstrated in Section 5, in the numerical experiments considered, it is usually sufficient to consider $M$ of the order of a few thousand to obtain results that are very stable and close to those that would be obtained with $M=2^{d}$. This observation is very useful when considering high dimensions (indicatively $d>10$) as it allows reducing the computational time that would otherwise tend to explode with dimension $d$. Finally, it is important to note that at each time step $t_{n}$, the calculation of the solutions to problems (4.3) for each value $\mathbf{x}^{n,p}\in X^{n}$ are independent of one another. Therefore, these problems can be addressed using parallel computing, which significantly reduces the overall computational time.
Remark 3.
The $\Gamma^{\text{avg}}$ matrix defined in (4.7), which is used to defined the sets $X_{n}$ for $n=1,\dots,N,$ may not be positive semidefinite. In this particular case, certain techniques can be employed to determine the closest positive semidefinite matrix to $\Gamma^{\text{avg}}$ according to the spectral norm. One of these techniques is called Projection onto the Cone of Positive Semidefinite Matrices (see e.g. Calafiore and El Ghaoui [3]). For the record, in the cases we analyzed in the numerical experiments Section (5), this situation never occurred.
5 Numerical Experiments
In this Section we report some numerical results in order to investigate the effectiveness of the proposed GTU algorithm for pricing European options in high dimension.
Numerical tests are divided into 3 categories. In the first group of tests, we assume that there is no uncertainty about correlation. In the second battery of tests the correlation between the various underlyings is uncertain. In the third set of tests we instead consider a path dependent option written on a single underlying asset. While in the first two cases, the high dimensionality of the problem arises because the option payoff depends on a large number of underlying assets, in the last case, the high dimensionality originates because the payoff depends on the values assumed by the single underlying asset at numerous previous times.
Parameters employed for the tests are shown in Table 1. The options chosen for these tests, along with their numerical parameters, are identical to those analyzed by Guyon and Labordère [6]. This allows us to directly compare our results with theirs.
The algorithm has been implemented in MATLAB and computations have been preformed on a server which employs a 2.40 GHz Intel Xenon processor (Gold 6148, Skylake) and 20 GB of RAM. To better evaluate the computational time of the algorithm, parallelization was not used for solving the optimization problems (4.3).
Model parameters Algorithm parameters
Symbol Meaning Value Symbol Meaning Value
$S_{0}^{i}$ Initial value of $S^{i}$ $100$ $K_{1}$ Strike of long call $90$
$\sigma_{i}^{\min}$ Lower bound for $\sigma_{i}$ $0.1$ $K_{2}$ Strike of short call $110$
$\sigma_{i}^{\max}$ Upper bound for $\sigma_{i}$ $0.2$ $K$ Strike of Call Sharpe $100$
$\rho_{i,j}^{\min}$ Lower bound for $\rho_{i,j}$ $-0.5$ $d$ Number of underlyings variable
$\rho_{i,j}^{\max}$ Upper bound for $\rho_{i,j}$ $+0.5$ $N$ Number of time steps variable
$r$ Risk free i.r. 00 $P$ Number of points in $X_{n}$ variable
$T$ Maturity $1.0$ $M$ Number of branches for GPR-Tree variable
Table 1: Parameters employed for the numerical tests.
5.1 Model with no uncertainty on correlation
In this first battery of tests, we assume that the correlation between the various Brownians is known, which is equivalent to imposing $\rho_{i,j}^{\min}=\rho_{i,j}^{\max}$. In particular, it can be proven that if all correlations are equal and non-negative, then the correlation matrix is positive semidefinite, ensuring that this parameter set is well-defined. The level of correlation adopted in the various tests is specified in the following tables.
5.1.1 Outperformer option
The Outperformer option is a financial derivative that pays a payoff equal to the positive part of the difference between two asset values at maturity. Specifically, the payoff is given by
$\Psi\left(\mathbf{S}_{T}\right)=\left(S_{T}^{2}-S_{T}^{1}\right)_{+}.$ (5.1)
Thus, the valuation of this option is a 2-dimensional problem, which can be further simplified to a problem in dimension 1 when $\rho_{1,2}\leq 0$ by employing the change of numèraire technique as described by Guyon and Labordère [6]. Thus, although the dimension of this problem is not large, it is interesting in that we have a reliable benchmark (computed for example with the PDE method by Windcliff et al. [13]) as well as the results reported in the paper by Guyon and Labordère [6].
Table 2 provides the numerical results. The results obtained with GTU closely match the benchmark, even for small values of $N$ and $P$. The computational time is minimal.
$N=16$ $N=32$ $N=64$
$\rho_{1,2}$ $P$ $125$ $250$ $500$ $1000$ $125$ $250$ $500$ $1000$ $125$ $250$ $500$ $1000$ $GL$ $BM$
$-0.5$ $\underset{\left(27\right)}{13.80}$ $\underset{\left(43\right)}{13.81}$ $\underset{\left(95\right)}{13.80}$ $\underset{\left(239\right)}{13.80}$ $\underset{\left(43\right)}{13.78}$ $\underset{\left(85\right)}{13.79}$ $\underset{\left(182\right)}{13.78}$ $\underset{\left(458\right)}{13.78}$ $\underset{\left(83\right)}{13.77}$ $\underset{\left(166\right)}{13.78}$ $\underset{\left(359\right)}{13.77}$ $\underset{\left(920\right)}{13.77}$ $13.77$ $13.75$
00 $\underset{\left(25\right)}{11.28}$ $\underset{\left(42\right)}{11.27}$ $\underset{\left(90\right)}{11.26}$ $\underset{\left(221\right)}{11.28}$ $\underset{\left(43\right)}{11.27}$ $\underset{\left(84\right)}{11.26}$ $\underset{\left(181\right)}{11.25}$ $\underset{\left(470\right)}{11.27}$ $\underset{\left(84\right)}{11.28}$ $\underset{\left(170\right)}{11.26}$ $\underset{\left(361\right)}{11.25}$ $\underset{\left(1022\right)}{11.27}$ $11.26$ $11.25$
$0.5$ $\underset{\left(25\right)}{8.03}$ $\underset{\left(44\right)}{7.99}$ $\underset{\left(87\right)}{7.99}$ $\underset{\left(225\right)}{8.00}$ $\underset{\left(43\right)}{8.04}$ $\underset{\left(83\right)}{7.99}$ $\underset{\left(191\right)}{7.98}$ $\underset{\left(457\right)}{7.99}$ $\underset{\left(83\right)}{8.07}$ $\underset{\left(167\right)}{8.01}$ $\underset{\left(377\right)}{7.99}$ $\underset{\left(1012\right)}{7.99}$
Table 2: Outperformer option. The column labeled “GL” presents the results from [6] for the same option, while the column labeled “BM” shows the benchmark computed by using the change of numéraire technique. Values in brackets indicate the computational times, measured in seconds.
5.1.2 Outperformer spread option
Similar to the Outperformer option discussed in Subsection 5.1.1, the Outperformer spread option is another example of 2-dimensional option which has been investigated by Guyon and Labordère [6] and for which, when $\rho_{1,2}\leq 0$, a benchmark is available by using the change of numèraire technique. Specifically, the payoff of such an option reads out
$\Psi\left(\mathbf{S}_{T}\right)=\left(S_{T}^{2}-0.9S_{T}^{1}\right)_{+}-\left(S_{T}^{2}-1.1S_{T}^{1}\right)_{+}.$ (5.2)
Results are presented in Table (3) are very close to the benchmark.
$N=32$ $N=64$ $N=128$
$\rho$ $P$ $125$ $250$ $500$ $1000$ $2000$ $125$ $250$ $500$ $1000$ $2000$ $125$ $250$ $500$ $1000$ $2000$ $GL$ $BM$
$-0.5$ $\underset{\left(42\right)}{11.31}$ $\underset{\left(82\right)}{11.32}$ $\underset{\left(176\right)}{11.34}$ $\underset{\left(463\right)}{11.34}$ $\underset{\left(1312\right)}{11.34}$ $\underset{\left(81\right)}{11.32}$ $\underset{\left(180\right)}{11.34}$ $\underset{\left(364\right)}{11.37}$ $\underset{\left(935\right)}{11.37}$ $\underset{\left(2642\right)}{11.37}$ $\underset{\left(170\right)}{11.33}$ $\underset{\left(356\right)}{11.35}$ $\underset{\left(778\right)}{11.38}$ $\underset{\left(2197\right)}{11.38}$ $\underset{\left(5442\right)}{11.39}$ $11.37$ $11.41$
00 $\underset{\left(42\right)}{11.29}$ $\underset{\left(81\right)}{11.32}$ $\underset{\left(193\right)}{11.31}$ $\underset{\left(532\right)}{11.32}$ $\underset{\left(1233\right)}{11.32}$ $\underset{\left(81\right)}{11.31}$ $\underset{\left(182\right)}{11.35}$ $\underset{\left(358\right)}{11.34}$ $\underset{\left(1036\right)}{11.35}$ $\underset{\left(2280\right)}{11.35}$ $\underset{\left(173\right)}{11.29}$ $\underset{\left(342\right)}{11.35}$ $\underset{\left(733\right)}{11.35}$ $\underset{\left(1981\right)}{11.36}$ $\underset{\left(5403\right)}{11.36}$ $11.39$
$0.5$ $\underset{\left(42\right)}{11.15}$ $\underset{\left(81\right)}{11.19}$ $\underset{\left(170\right)}{11.19}$ $\underset{\left(446\right)}{11.18}$ $\underset{\left(1078\right)}{11.18}$ $\underset{\left(80\right)}{11.14}$ $\underset{\left(186\right)}{11.20}$ $\underset{\left(361\right)}{11.20}$ $\underset{\left(904\right)}{11.20}$ $\underset{\left(2309\right)}{11.19}$ $\underset{\left(165\right)}{11.13}$ $\underset{\left(339\right)}{11.20}$ $\underset{\left(781\right)}{11.20}$ $\underset{\left(2038\right)}{11.20}$ $\underset{\left(5081\right)}{11.20}$
Table 3: Outperformer spread option. The column labeled “GL” presents the results from [6] for the same option, while the column labeled “BM” shows the benchmark computed by using the change of numéraire technique. Values in brackets indicate the computational times, measured in seconds.
5.1.3 Geo-Call spread
The Geo-Call spread option generalized the Call spread option discussed in Subsection 5.1.2 to a multi-asset option. Specifically, the payoff is given by the difference between the payoffs of two call options written on the geometric mean of the underlyings considered.
$\Psi\left(\mathbf{S}_{T}\right)=\left(\sqrt[d]{\prod_{k=1}^{d}S_{T}^{k}}-K_{1}\right)_{+}-\left(\sqrt[d]{\prod_{k=1}^{d}S_{T}^{k}}-K_{2}\right)_{+}$
The use of the geometric mean is of particular interest since, if the underlyings follow (2.5) and $\rho_{i,j}=0$, than this expression has the same probability distribution of a single underlying as shown below:
$d]{\prod_{k=1}^{d}S_{T}^{k}}\sim\sqrt[d]{\prod_{k=1}^{d}S_{0}^{k}}\exp\left(\left(r-\hat{\eta}-\frac{1}{2}\left(\frac{1}{d}\sqrt{\sum_{k=1}^{d}\sigma_{i}^{2}}\right)^{2}\right)T+\left(\frac{1}{d}\sqrt{\sum_{k=1}^{d}\sigma_{i}^{2}}\right)B_{T}\right)$
with $B$ a Brownian motion and $\hat{\eta}$ the modified dividend yield, defined as
$\hat{\eta}=\frac{1}{d}\sum_{k=1}^{d}\eta_{i}+\frac{d-1}{2d^{2}}\sum_{k=1}^{d}\sigma_{i}^{2}.$
Therefore, a problem can be traced back to dimension 1 in which the volatility
$\hat{\sigma}=\frac{1}{d}\sqrt{\sum_{k=1}^{d}\sigma_{i}^{2}}$
ranges between
$\frac{1}{d}\sqrt{\sum_{k=1}^{d}\left(\sigma_{i}^{\min}\right)^{2}}=\min_{\sigma_{i}\in\left[\sigma_{i}^{\min},\sigma_{i}^{\max}\right]}\left(\frac{1}{d}\sqrt{\sum_{k=1}^{d}\sigma_{i}^{2}}\right)\leq\hat{\sigma}\leq\max_{\sigma_{i}\in\left[\sigma_{i}^{\min},\sigma_{i}^{\max}\right]}\left(\frac{1}{d}\sqrt{\sum_{k=1}^{d}\sigma_{i}^{2}}\right)=\frac{1}{d}\sqrt{\sum_{k=1}^{d}\left(\sigma_{i}^{\max}\right)^{2}}$
and the dividend $\hat{\eta}$ is equal to
$\hat{\eta}=\frac{1}{d}\sum_{k=1}^{d}\eta_{i}+\frac{d-1}{2d^{2}}\sum_{k=1}^{d}\sigma_{i}^{2}=\frac{1}{d}\sum_{k=1}^{d}\eta_{i}+\frac{d-1}{2}\hat{\sigma}^{2}.$
This case is a generalization of the UVM model in that the dividend is not constant but changes in relation to the value of volatility. The binomial tree method proposed by Avellaneda et al. [1] for the 1d UVM model can be used to calculate the price in this model, taking care to change the dividend in relation to volatility. In this regard, it should be noted that while in the one-dimensional UVM model the optimal $\sigma$ value is always $\sigma^{\min}$ or $\sigma^{\max}$ (this feature is known as “the bang bang condition”), in this model, characterized by a variable dividend yield $\hat{\eta}$, this property no longer applies. Therefore, at each node of the tree, the choice of optimal correlation is to be pursued not merely by comparing the objective function at the extreme values of $\sigma$, but by optimising the option price via an appropriate algorithm, such as the Sequential Quadratic Programming algorithm.
The results are shown in Table 4. As shown, the values obtained for varying $N$ and $P$ are very close to the benchmark for the considered values of $d$. The largest deviation is observed for $d=5$, with a difference of $0.04$ between the GTU price and the benchmark. However, this difference is minor, being less than $0.5\%$ of the benchmark. Additionally, we observe that the computational time increases linearly with the number of time steps $N$ and approximately linearly with the number of points $P$. This is particularly interesting since the cost for the GPR regression is cubic in $P$, indicating that most of the computational time is spent on solving optimization problems rather than on the regression itself. Lastly, the computational time grows exponentially with problem size, as the number of values to be evaluated with a tree step is $2^{d}$. Despite this, we are able to handle high dimensions, such as $d=10$, with reasonable computational effort.
$N=16$ $N=32$ $N=64$ $BM$
$d$ $P$ $125$ $250$ $500$ $1000$ $125$ $250$ $500$ $1000$ $125$ $250$ $500$ $1000$
$2$ $\underset{\left(25\right)}{10.48}$ $\underset{\left(48\right)}{10.47}$ $\underset{\left(104\right)}{10.47}$ $\underset{\left(277\right)}{10.47}$ $\underset{\left(45\right)}{10.49}$ $\underset{\left(84\right)}{10.48}$ $\underset{\left(192\right)}{10.49}$ $\underset{\left(440\right)}{10.48}$ $\underset{\left(92\right)}{10.49}$ $\underset{\left(177\right)}{10.49}$ $\underset{\left(414\right)}{10.49}$ $\underset{\left(950\right)}{10.49}$ $10.50$
$5$ $\underset{\left(33\right)}{9.76}$ $\underset{\left(62\right)}{9.74}$ $\underset{\left(139\right)}{9.70}$ $\underset{\left(380\right)}{9.70}$ $\underset{\left(68\right)}{9.77}$ $\underset{\left(129\right)}{9.77}$ $\underset{\left(274\right)}{9.72}$ $\underset{\left(830\right)}{9.72}$ $\underset{\left(129\right)}{9.76}$ $\underset{\left(286\right)}{9.79}$ $\underset{\left(698\right)}{9.73}$ $\underset{\left(1945\right)}{9.74}$ $9.70$
$10$ $\underset{\left(188\right)}{9.45}$ $\underset{\left(590\right)}{9.50}$ $\underset{\left(2119\right)}{9.51}$ $\underset{\left(6900\right)}{9.54}$ $\underset{\left(468\right)}{9.45}$ $\underset{\left(1519\right)}{9.49}$ $\underset{\left(5582\right)}{9.51}$ $\underset{\left(18892\right)}{9.54}$ $\underset{\left(1171\right)}{9.44}$ $\underset{\left(3738\right)}{9.48}$ $\underset{\left(13904\right)}{9.51}$ $\underset{\left(50299\right)}{9.52}$ $9.55$
Table 4: Geo-Call spread, with $\rho_{i,j}=0$ and with $M=2^{d}$. Values in brackets represent the computational times, measured in seconds.
To prevent the computational time from increasing exponentially, we can reduce the number of terms in the sum by randomly selecting a sample of $M$ elements. This procedure is justifiable in that, as $M$ approaches $2^{d}$ , the calculated price converges to the value we would get without this simplification. The results, shown in Table 5, indicate that even with just a few thousand samples, the outcomes are very stable. In addition, we can achieve high size of $d=40$ with reasonable computation time.
$N=16$ $N=32$ $N=64$
$d$ $M$ $P$ $125$ $250$ $500$ $1000$ $125$ $250$ $500$ $1000$ $125$ $250$ $500$ $1000$ $BM$
[…]
$1000$ $\underset{\left(197\right)}{9.45}$ $\underset{\left(589\right)}{9.50}$ $\underset{\left(2006\right)}{9.51}$ $\underset{\left(6868\right)}{9.54}$ $\underset{\left(465\right)}{9.45}$ $\underset{\left(1520\right)}{9.49}$ $\underset{\left(5461\right)}{9.51}$ $\underset{\left(18118\right)}{9.54}$ $\underset{\left(1360\right)}{9.44}$ $\underset{\left(3856\right)}{9.48}$ $\underset{\left(15310\right)}{9.51}$ $\underset{\left(50299\right)}{9.52}$
[…]
$1000$ $\underset{\left(175\right)}{9.44}$ $\underset{\left(619\right)}{9.51}$ $\underset{\left(2422\right)}{9.56}$ $\underset{\left(8330\right)}{9.57}$ $\underset{\left(513\right)}{9.43}$ $\underset{\left(1662\right)}{9.51}$ $\underset{\left(6004\right)}{9.55}$ $\underset{\left(22278\right)}{9.56}$ $\underset{\left(1513\right)}{9.43}$ $\underset{\left(4525\right)}{9.51}$ $\underset{\left(16695\right)}{9.55}$ $\underset{\left(60102\right)}{9.55}$
[…]
$1000$ $\underset{\left(389\right)}{9.33}$ $\underset{\left(1392\right)}{9.40}$ $\underset{\left(4396\right)}{9.48}$ $\underset{\left(16209\right)}{9.53}$ $\underset{\left(1058\right)}{9.32}$ $\underset{\left(3997\right)}{9.39}$ $\underset{\left(13458\right)}{9.49}$ $\underset{\left(46350\right)}{9.54}$ $\underset{\left(3484\right)}{9.30}$ $\underset{\left(11411\right)}{9.39}$ $\underset{\left(36140\right)}{9.49}$ $\underset{\left(129875\right)}{9.53}$
Table 5: Geo-Call spread in high dimension, for $\rho_{i,j}=0$ and $M<2^{d}$ variable. Values in brackets represent the computational times, measured in seconds.
The results shown in the Tables 4 and 5 are for unrelated underlying assets. To demonstrate that the GTU algorithm also works well with correlated underlyings, we assumed all correlations equal to $0.50$ or $0.75$. The results are shown in Table 6. In this case, we do not have a benchmark for comparison. However, the obtained values demonstrate a high degree of stability, suggesting their accuracy.
$N=16$ $N=32$ $N=64$
$d$ $P$ $125$ $250$ $500$ $1000$ $125$ $250$ $500$ $1000$ $125$ $250$ $500$ $1000$
CASE $\rho_{i,j}=0.50$
$2$ $\underset{\left(26\right)}{10.90}$ $\underset{\left(43\right)}{10.89}$ $\underset{\left(95\right)}{10.88}$ $\underset{\left(211\right)}{10.88}$ $\underset{\left(46\right)}{10.92}$ $\underset{\left(91\right)}{10.91}$ $\underset{\left(181\right)}{10.91}$ $\underset{\left(459\right)}{10.90}$ $\underset{\left(90\right)}{10.93}$ $\underset{\left(178\right)}{10.92}$ $\underset{\left(376\right)}{10.91}$ $\underset{\left(910\right)}{10.91}$
$5$ $\underset{\left(30\right)}{10.71}$ $\underset{\left(57\right)}{10.68}$ $\underset{\left(131\right)}{10.68}$ $\underset{\left(343\right)}{10.66}$ $\underset{\left(52\right)}{10.75}$ $\underset{\left(115\right)}{10.73}$ $\underset{\left(259\right)}{10.72}$ $\underset{\left(762\right)}{10.70}$ $\underset{\left(107\right)}{10.75}$ $\underset{\left(240\right)}{10.75}$ $\underset{\left(588\right)}{10.74}$ $\underset{\left(1688\right)}{10.72}$
$10$ $\underset{\left(150\right)}{10.62}$ $\underset{\left(532\right)}{10.66}$ $\underset{\left(1748\right)}{10.62}$ $\underset{\left(6214\right)}{10.59}$ $\underset{\left(398\right)}{10.64}$ $\underset{\left(1234\right)}{10.70}$ $\underset{\left(4284\right)}{10.67}$ $\underset{\left(16489\right)}{10.64}$ $\underset{\left(906\right)}{10.62}$ $\underset{\left(2895\right)}{10.72}$ $\underset{\left(10497\right)}{10.69}$ $\underset{\left(38313\right)}{10.68}$
CASE $\rho_{i,j}=0.75$
$2$ $\underset{\left(27\right)}{11.02}$ $\underset{\left(46\right)}{11.05}$ $\underset{\left(84\right)}{11.05}$ $\underset{\left(212\right)}{11.04}$ $\underset{\left(44\right)}{11.03}$ $\underset{\left(82\right)}{11.06}$ $\underset{\left(175\right)}{11.06}$ $\underset{\left(441\right)}{11.06}$ $\underset{\left(83\right)}{11.03}$ $\underset{\left(163\right)}{11.06}$ $\underset{\left(350\right)}{11.07}$ $\underset{\left(909\right)}{11.07}$
$5$ $\underset{\left(28\right)}{11.00}$ $\underset{\left(51\right)}{10.98}$ $\underset{\left(115\right)}{10.96}$ $\underset{\left(347\right)}{10.96}$ $\underset{\left(55\right)}{11.03}$ $\underset{\left(116\right)}{11.03}$ $\underset{\left(249\right)}{11.00}$ $\underset{\left(808\right)}{11.00}$ $\underset{\left(106\right)}{11.04}$ $\underset{\left(234\right)}{11.06}$ $\underset{\left(536\right)}{11.02}$ $\underset{\left(1839\right)}{11.03}$
$10$ $\underset{\left(129\right)}{11.07}$ $\underset{\left(491\right)}{11.03}$ $\underset{\left(1695\right)}{10.98}$ $\underset{\left(6714\right)}{10.96}$ $\underset{\left(378\right)}{11.10}$ $\underset{\left(1108\right)}{11.08}$ $\underset{\left(4126\right)}{11.03}$ $\underset{\left(15925\right)}{11.02}$ $\underset{\left(816\right)}{11.09}$ $\underset{\left(2695\right)}{11.10}$ $\underset{\left(10790\right)}{11.05}$ $\underset{\left(38559\right)}{11.05}$
Table 6: Geo-Call spread with correlated underlyings. Values in brackets represent the computational times, measured in seconds.
5.2 Model with uncertainty on correlation
We now test the GTU algorithm under the assumption that the correlation coefficients $\rho_{i,j}$ are also unknown, let them vary between two bounds $\rho_{i,j}^{\min}=-0.5$ and $\rho_{i,j}^{\max}=0.5$. Evaluating an option in this scenario is more complex for two reasons. Firstly, it is necessary to verify that the correlation matrix $\Gamma$, resulting from varying the coefficients $\rho_{i,j}$ during the optimization procedure, remains positive semidefinite. This verification is implemented directly by using MATLAB’s chol function. Secondly, the number of these coefficients increases quadratically with dimension, leading to a rapid increase in computational cost, making it challenging to obtain results in high dimensions.
5.2.1 Outperformer spread option with variable correlation
The payoff of this option is given in (5.2). We only consider the case $d=2$ since this is the only case considered by Guyon and Labordère [6]. Moreover, in this particular case it is unnecessary to verify that the correlation matrix $\Gamma$ is positive semidefinite, as this holds true for any permissible value of $\rho_{1,2}$. We propose this particular case because it has also been studied by Guyon and Labordère [6] and thus it is possible to test our algorithm against theirs. In this regard, we can observe that our results reported in Table 7 are slightly higher than the reference value in [6]. This slight difference probably results from the fact that our optimization procedure performs better than Guyon and Labordère’s one in [6] and therefore, it is better able to approach the optimal price in UVM. We also note that the price obtained with a variable correlation parameter, specifically $-0.5\leq\rho_{1,2}\leq 0.5$, is significantly higher than the prices obtained considering a fixed correlation of $\rho_{1,2}=0.5$, $\rho_{1,2}=0$ or $\rho_{1,2}=-0.5$, shown in Table 3. Specifically, the higher result for a fixed correlation is achieved for $\rho_{1,2}=-0.5$ and it is equal to $11.39$, while for a variable correlation, we reach $12.77$. To conclude, it is crucial to emphasize the importance of considering a robust model for correlation among the underlying assets, as pricing results can vary significantly. Therefore, variable correlation is a risk factor that practitioners should not overlook.
$N/P$ $125$ $250$ $500$ $1000$ $2000$ $GL$
$32$ $\underset{\left(129\right)}{12.60}$ $\underset{\left(242\right)}{12.72}$ $\underset{\left(577\right)}{12.72}$ $\underset{\left(1754\right)}{12.73}$ $\underset{\left(2302\right)}{12.74}$ $12.67$
$64$ $\underset{\left(380\right)}{12.59}$ $\underset{\left(537\right)}{12.73}$ $\underset{\left(1244\right)}{12.74}$ $\underset{\left(3490\right)}{12.76}$ $\underset{\left(4791\right)}{12.77}$
$128$ $\underset{\left(355\right)}{12.52}$ $\underset{\left(733\right)}{12.69}$ $\underset{\left(1466\right)}{12.72}$ $\underset{\left(4116\right)}{12.75}$ $\underset{\left(10331\right)}{12.77}$
Table 7: Outperformer spread option with variable correlation $\rho_{1,2}$. The column labeled “GL” presents the results from [6] for the same option. Values in brackets represent the computational times, measured in seconds.
5.2.2 Geo-Outperformer option with variable correlation
In Subsection 5.1.1 we have introduced the Outperformer option, in which the payoff (5.1) depends only on 2 underlyings. In this Subsection, we extend the definition of such an option to an option on multiple underlyings. Specifically, we replace the second underlying with the geometric mean of all the underlying with the exception of the first one, so the payoff function reads out:
$\Psi\left(\mathbf{S}_{T}\right)=\left(\sqrt[d-1]{\prod_{k=2}^{d}S_{T}^{k}}-S_{T}^{1}\right)_{+}$ (5.3)
In this particular case, the optimal combination of the $\rho_{i,j}$ coefficients is easily found: maximize the correlation between all the underlyings except the first one and minimize the correlation between the first underlying and the remaining ones. In our numerical setting, this implies
$\rho_{i,j}=\begin{cases}1&\text{if }i=j\\ -0.5&\text{if i$\neq j$ and}\\ 0.5&\text{otherwise.}\end{cases}\min\left(i,j\right)=1$ (5.4)
This solutions, in fact, maximizes the volatility of the difference in (5.3) and thus the expected value of the payoff. This remark allows us to compute a benchmark by means of the GTU algorithm for correlation fixed according to (5.4). In particular, we compute the Benchmark by using $N=128$ time steps and $P=1000$ points.
The results presented in Table 8 are generally consistent with the benchmark, albeit slightly lower. This discrepancy is likely due to the method used to construct the sets $X_{n}$ required for GPR. Specifically, as far as the benchmark is considered, these sets are constructed using optimal correlation in (5.4), whereas for the general problem with variable correlation, they are determined by using the average correlation (4.8), which is sub-optimal. Consequently, more points are needed, compared to the fixed correlation model, to achieve results of equivalent quality.
$N=32$ $N=64$ $N=128$ $BM$
$d$ $P$ $125$ $250$ $500$ $1000$ $2000$ $125$ $250$ $500$ $1000$ $2000$ $125$ $250$ $500$ $1000$ $2000$
$2$ $\underset{\left(70\right)}{13.78}$ $\underset{\left(138\right)}{13.79}$ $\underset{\left(304\right)}{13.78}$ $\underset{\left(941\right)}{13.78}$ $\underset{\left(2491\right)}{13.78}$ $\underset{\left(146\right)}{13.76}$ $\underset{\left(301\right)}{13.78}$ $\underset{\left(719\right)}{13.77}$ $\underset{\left(2127\right)}{13.76}$ $\underset{\left(6283\right)}{13.76}$ $\underset{\left(312\right)}{13.74}$ $\underset{\left(671\right)}{13.78}$ $\underset{\left(1548\right)}{13.77}$ $\underset{\left(4840\right)}{13.76}$ $\underset{\left(12463\right)}{13.76}$ $13.75$
$3$ $\underset{\left(168\right)}{12.84}$ $\underset{\left(368\right)}{13.00}$ $\underset{\left(758\right)}{12.98}$ $\underset{\left(1982\right)}{12.96}$ $\underset{\left(4335\right)}{12.95}$ $\underset{\left(421\right)}{12.82}$ $\underset{\left(768\right)}{12.99}$ $\underset{\left(1554\right)}{12.96}$ $\underset{\left(4350\right)}{12.94}$ $\underset{\left(9336\right)}{12.93}$ $\underset{\left(900\right)}{12.78}$ $\underset{\left(1633\right)}{12.97}$ $\underset{\left(3154\right)}{12.96}$ $\underset{\left(8221\right)}{12.94}$ $\underset{\left(17646\right)}{12.93}$ $12.96$
$4$ $\underset{\left(384\right)}{12.46}$ $\underset{\left(716\right)}{12.64}$ $\underset{\left(1165\right)}{12.67}$ $\underset{\left(2632\right)}{12.71}$ $\underset{\left(7216\right)}{12.70}$ $\underset{\left(820\right)}{12.37}$ $\underset{\left(1349\right)}{12.60}$ $\underset{\left(2086\right)}{12.64}$ $\underset{\left(5156\right)}{12.70}$ $\underset{\left(11520\right)}{12.69}$ $\underset{\left(1833\right)}{12.25}$ $\underset{\left(2630\right)}{12.53}$ $\underset{\left(4083\right)}{12.60}$ $\underset{\left(10019\right)}{12.69}$ $\underset{\left(22713\right)}{12.69}$ $12.73$
$5$ $\underset{\left(2006\right)}{12.15}$ $\underset{\left(3431\right)}{12.52}$ $\underset{\left(2022\right)}{12.53}$ $\underset{\left(6328\right)}{12.62}$ $\underset{\left(9796\right)}{12.57}$ $\underset{\left(1847\right)}{12.05}$ $\underset{\left(2646\right)}{12.45}$ $\underset{\left(3918\right)}{12.49}$ $\underset{\left(10362\right)}{12.60}$ $\underset{\left(18192\right)}{12.56}$ $\underset{\left(3534\right)}{11.90}$ $\underset{\left(5103\right)}{12.35}$ $\underset{\left(7420\right)}{12.41}$ $\underset{\left(20298\right)}{12.57}$ $\underset{\left(54465\right)}{12.54}$ $12.64$
Table 8: Outperformer geo option with variable correlation. Values in brackets represent the computational times, measured in seconds. The column labeled “BM” shows the benchmark computed by using the GTU algorithm with the optimal correlation parameters.
5.3 Path dependent option: the Call Sharpe option
We conclude our numerical analysis by discussing an option known as Call Sharpe, previously analyzed by Guyon and Labordére in [6]. This option is particularly interesting as it exemplifies a path-dependent option, where the value depends not only on the current price of the underlying asset but also on the previously realized volatility. Specifically, the payoff at maturity, which depends on a single underlying asset, and on its realized volatility which is computed on a monthly base. To this aim, we assume that the maturity time $T$ is a multiple of $1/12$, defining $N_{m}=12T\in\mathbb{N}$ as the number of months comprising the life of the contract. The realized volatility computed using monthly returns is given by
$V_{T}=\frac{1}{T}\sum_{l=1}^{N_{m}}\left(\ln\left(\frac{S_{\tau_{l}}}{S_{\tau_{l-1}}}\right)\right)^{2},$
with $\tau_{l}=l/12$, and the payoff at maturity is
$\Psi\left(S_{T}\right)=\frac{\left(S_{T}-K\right)^{+}}{\sqrt{V_{T}}}.$
Following Guyon and Labordére [6], at time $t$, the option value depends on $S_{t}$ and on the two path-dependent variables
$A_{t}^{1}$ =$=\sum_{\{l\mid\tau_{l}\leq t\}}\left(\ln\left(\frac{S_{\tau_{l}}}{S_{\tau_{l-1}}}\right)\right)^{2},$
$A_{t}^{2}$ $=S_{\sup_{\{l\mid\tau_{l}\leq t\}}\tau_{l}}.$
Moreover, we stress out that, if $t=\tau_{l}$ for a certain $l$ value, than $A_{t}^{2}=S_{t}$, so we can discard this variate when $t$ is a multiple of $1/12$, that is $12t\in\mathbb{N}$.
Since this option is path dependent, some adaptations are required to use GTU. Let us begin by pointing out that the number of time steps $N$ must be a multiple of $N_{m}$, so that the monitoring dates $\tau_{l}$ are included in the time grid.
Next, we note that the elements of the sets $X_{n}$, used to define the points at which to evaluate the contract value, are determined through Monte Carlo simulations. Specifically, first we simulate $P$ randoms paths of the underlying stock price:
$\left\{S_{t_{n}}^{p}=S_{t_{n-1}}^{p}e^{\left(r-\eta-\left(\sigma^{\text{avg}}\right)^{2}/2\right)\Delta t+\sigma^{\text{avg}}\sqrt{\Delta t}Z^{n,p}},n=1,\dots,N,p=1,\dots,P\right\},$ (5.5)
with $S_{0}^{p}=S_{0}$, $\Delta t=T/N$ and $Z^{n,p}\sim\mathcal{N}\left(0,1\right)$. Then, the elements $\mathbf{x}^{n,p}$ of $X_{n}$ are defined as
$\mathbf{x}^{n,p}=\begin{cases}\left(S_{t_{n}}^{p},A_{t_{n}}^{1,p}\right)&\text{if }12t_{n}\in\mathbb{N},\\ \left(S_{t_{n}}^{p},A_{t_{n}}^{1,p},A_{t_{n}}^{2,p}\right)&\text{otherwise},\end{cases}$
with
$A_{t_{n}}^{1,p}=\sum_{\{l\mid\tau_{l}\leq t_{n}\}}\left(\ln\left(\frac{S_{\tau_{l}}^{p}}{S_{\tau_{l-1}}^{p}}\right)\right)^{2},\quad A_{t_{n}}^{2,p}=S_{\sup_{\{l\mid\tau_{l}\leq t\}}\tau_{l}}^{p}.$
We stress out that, as far as this path dependent option is considered, we opt for the Monte Carlo method over the quasi-Monte Carlo method to generate the points $\mathbf{x}^{n,p}$ of $X_{n}$ in (5.5), as the dimension of random simulations equals $N_{m}$, to much to make quasi-Monte Carlo be effective with only a few hundreds of samples.
For a particular value $\hat{\sigma}$ for the volatility, determined during the optimization procedure, the future points for $\mathbf{x}^{n,p}$ are only 2, namely $\tilde{\mathbf{x}}_{i}^{n,p,1}$ and $\tilde{\mathbf{x}}_{i}^{n,p,2}$, defined as
$\tilde{\mathbf{x}}_{i}^{n,p,m}=\begin{cases}\left(S_{t_{n}}^{p,m},A_{t_{n}}^{1,p}+\left(\ln\left(\frac{S_{t_{n}}^{p,m}}{A_{t_{n}}^{2}}\right)\right)^{2}\right)&\text{if }12t_{n+1}\in\mathbb{N},\\ \left(S_{t_{n}}^{p,m},A_{t_{n}}^{1,p}+\left(\ln\left(\frac{S_{t_{n}}^{p,m}}{A_{t_{n}}^{2}}\right)\right)^{2},A_{t_{n}}^{2,p}\right)&\text{otherwise}.\end{cases}$
for $m=1,2$ and with
$S_{t_{n}}^{p,m}=S_{t_{n}}^{p}e^{\left(r-\eta-\left(\hat{\sigma}\right)^{2}/2\right)\Delta t+\hat{\sigma}\sqrt{\Delta t}\left(-1\right)^{m}}.$
Moreover, as far as this path dependent option is considered, the elements that identify the market state, namely $S,A^{1}$ and $A^{2}$, represent very different quantities from each other. For this reason, as remarked in a similar framework by Goudènege et al. [5], it is beneficial to use a kernel that considers a different length scale for each predictor, such as the ARD Matern 3/2 kernel $k_{\text{M3/2}}^{\text{ARD}}$ which is defines as follows:
$k_{\text{M3/2}}^{\text{ARD}}(\mathbf{x},\mathbf{x}^{\prime})=\sigma_{f}^{2}\left(1+\sqrt{3\sum\left(\frac{\mathbf{x}_{i}-\mathbf{x}^{\prime}_{i}}{\sigma_{l,i}}\right)^{2}}\right)\exp\left(-\frac{\sqrt{3}}{\sigma_{l}}\sqrt{3\sum\left(\frac{\mathbf{x}_{i}-\mathbf{x}^{\prime}_{i}}{\sigma_{l,i}}\right)^{2}}\right),$
where $\sigma_{l,i}$ is the length scale for the $i$-th predictor.
In the specific case study considered in our numerical test, we consider $T=1$, so $N_{m}=12$. Results are reported in Table (9). The benchmark value, computed through a PDE approach described in [6], is $58.4$. In comparison, our more reliable price – obtained for $N=384$ and $P=2000$ – yields a value of $57.93$, much closer to the benchmark than the best price by Guyon and Labordère [6], which is worth $55.55$. Additionally, we emphasize that the relative deviation between our price and the benchmark is less than 1%, indicating a very small error. To conclude, we observe that, in this case, a large number of $N$ time steps is necessary to achieve accurate results. This requirement for many time steps necessitates the use of an algorithm with low computational cost per time step. Our tree method, which in this case requires the evaluation of only two future nodes, proves to be an appropriate choice.
$N\backslash P$ $250$ $500$ $1000$ $2000$ $GL$ $BM$
$24$ $\underset{\left(59\right)}{55.97}$ $\underset{\left(109\right)}{55.94}$ $\underset{\left(373\right)}{56.04}$ $\underset{\left(1325\right)}{56.35}$ $55.55$ $58.4$
$48$ $\underset{\left(98\right)}{59.35}$ $\underset{\left(217\right)}{57.26}$ $\underset{\left(751\right)}{57.14}$ $\underset{\left(2030\right)}{57.32}$
$96$ $\underset{\left(199\right)}{57.74}$ $\underset{\left(556\right)}{57.71}$ $\underset{\left(1580\right)}{57.46}$ $\underset{\left(4494\right)}{57.66}$
$192$ $\underset{\left(386\right)}{59.83}$ $\underset{\left(1016\right)}{57.30}$ $\underset{\left(3233\right)}{58.19}$ $\underset{\left(10114\right)}{57.90}$
$384$ $\underset{\left(872\right)}{58.19}$ $\underset{\left(2566\right)}{58.93}$ $\underset{\left(7315\right)}{58.22}$ $\underset{\left(26228\right)}{57.93}$
Table 9: Call Sharpe. The column labeled “GL” presents the results from [6] for the same option, while the column labeled “BM” shows the benchmark computed by using a PDE approach. Values in brackets indicate the computational times, measured in seconds.
6 Conclusions
In this study, we demonstrated the efficacy of Machine Learning in enhancing the Uncertain Volatility Model for high-dimensional option pricing. Our findings suggest that Gaussian Process Regression algorithm can effectively handle the complexity and variability inherent in financial markets, providing more accurate and reliable pricing under the UVM framework. The integration of Machine Learning not only improves the precision of the optimal volatility and correlation estimates but also optimizes computational efficiency, making it a valuable tool for traders and risk managers. Furthermore, our approach offers a robust method for superhedging strategies, ensuring non-negative profit and loss outcomes even in highly volatile market conditions. Future research can extend this work by exploring other ML techniques and their applications in different financial models, further bridging the gap between theoretical advancements and practical implementations in the field of quantitative finance.